Solve by the elimination method.
0.3x-0.2y=4
0.2x+0.5y=-75/17

Found 2 solutions by mananth, ikleyn:
Answer by mananth(16949)

-3x-0.2y=4 .............1
0.2x+0.5 y=-4.41 .............2
Eliminate y
multiply (1)by 0.5
Multiply (2) by 0.2
-1.5x-0.1y=2
0.04x +0.1y=-0.88
Add the two equations
-1.46x = 1.12
/-1.46
x=-0.77
plug value of x in (1)
-3x-0.2y =4
2.3-0.2y =4
-0.2y=4 -2.3
-0.2y=1.7
y=-8.52
m.ananth@hotmail.ca

Answer by ikleyn(53570)

.
Solve by the elimination method.
0.3x-0.2y=4
0.2x+0.5y=-75/17
~~~~~~~~~~~~~~~~~~~~~~~~~~

The solution x = -0.77, y = -8.52 in the post by @mananth is INCORRECT.

I checked it by substituting his values into the first equation ang got the value 1.473,
different from 4, which is in the right side of this equation.

It is just enough to refute the @mananth solution.